# Unsupervised Cross-Domain 3D Human Pose Estimation via Pseudo-Label-Guided Global Transforms

### v1

@@ IV-D Ablation Study
Effect of method’s modules. We conduct an ablation study starting from a baseline model that was trained on the source domain, with our method’s modules systematically integrated to analyze their effects: Pose Augmentor $M_{a}$, Pseudo-Label Generation module $M_{p}$, Global Transformation module $M_{t}$, and iterative training.
As shown in Table V, when our Pose Augmentor $M_{a}$ is added we achieve an improvement of 6.1$\downarrow$ in MPJPE, followed by a substantial error reduction of 20.6$\downarrow$ when the Global Transformation module $M_{t}$ with pseudo-labels from $M_{p}$ are introduced. Finally, through an iteration of the whole process, a further drop of 0.8$\downarrow$ in MPJPE is achieved. Corresponding increases in PCK and AUC are also shown.
TABLE V: The impact of each module in proposed method. The * indicates that the modules were processed for an extra iteration.
@@
VideoPose3D [23] + $M_{a}$ + ($M_{p}$ $M_{t}$) 58.6 91.8 64.3
VideoPose3D [23] + $M_{a}$ + ($M_{p}$ $M_{t}$)* 57.7 92.5 64.6
Different Global Transformations. The global transformation module $M_{t}$ contributes significantly in reducing the domain differences in camera positions and orientations from source to target. It is target-oriented and uses the pseudo-labels to guide the rotation and translation of the 3D poses. We compare three other approaches to a global transformation: RR+RT, GR+RT, RR+GT, where RR denotes rotation matrices are randomly initialized and learned in the network, RT denotes the translation matrices are randomly initialized and learned in the network, GR denotes the pseudo-label guided rotation and GT denotes the pseudo-label guided translation. Our approach is therefore GR+GT, which performs best as shown in Table VI.
TABLE VI: The impact of different approaches to Global Transformation $M_{t}$.
@@
IV-E Discussion Compared to SOTA Methods
@@
The Dual-Augmentor framework[20] generates diverse out-of-source distributions using a weak and strong augmentor, but it does not exploit 2D poses from the target domain, limiting its performance in cross-dataset evaluation. Our method fully utilizes target-domain 2D poses to generate pseudo-3D poses, guiding global transformations that better align 3D across domains. Additionally, our Pose Discriminator incorporates 2D poses from the target domain, ensuring that generated 2D poses also align with the target distribution.
@@ V Conclusion
This paper proposes an iterative framework for domain-adaptive human pose estimation, focusing on aligning the 2D and 3D poses across domains to bridge domain gaps. A Pseudo-Label guided Global Transformation module incorporates a human-centered coordinate system to explicitly align the global coordinates between source and target domains. This design is both intuitive and effective, achieving significant performance gains without adding extra training parameters. Additionally, we enhance pose augmentation by modifying bone lengths and angles, and we propose to use an adversarial learning architecture based on the target domain’s 2D poses and the source domain’s 3D poses.
@@ Acknowledgments
Algorithm 1 Computation of the transformation matrices $\bm{R},\bm{T}$ from camera coordinate system $C$ to human-centric coordinate system $H$.
Input: coordinates of 3D pose ${\bm{P}}_{j}=(X_{j},Y_{j},Z_{j}),(j=0,...,J-1)$ under $C$, where $j$ denotes the $j$-th joint.
Output: rotation matrix $\bm{R}$ and translation matrix $\bm{T}$
// Computation of the unit vector along x-axis, which is pointing from the left hip (4th joint) to the right hip (1st joint)
$\bm{v}_{1\_4}=\bm{P}_{1}-\bm{P}_{4}$
$\bm{e}_{1}=\bm{v}_{1\_4}/\left\|\bm{v}_{1\_4}\right\|$
// Computation of the unit vector along z-axis, which is perpendicular to the plane comprising the right hip (1st joint), left hip (4th joint) and thorax (8th joint)
$\bm{v}_{0\_8}=\bm{P}_{8}-(\bm{P}_{1}+\bm{P}_{4})/2$
$\bm{k}_{t}=\bm{e}_{1}\times\bm{v}_{0\_8}$
$\bm{e}_{3}=\bm{k}_{t}/\left\|\bm{k}_{t}\right\|$
// Computation of the unit vector along y-axis, which is perpendicular to the x and z axes
$\bm{e}_{2}=\bm{e}_{1}\times\bm{e}_{3}$
$\bm{\tilde{R}}=[\bm{e}_{1};\bm{e}_{2};\bm{e}_{3}]$
$\bm{\tilde{T}}=(\bm{P}_{1}+\bm{P}_{4})/2$
$\bm{R}=\bm{\tilde{R}}^{-1}$
$\bm{T}=-\bm{\tilde{T}}$
@@
Zhiyong Wang (Member, IEEE) received the Ph.D. degree in mechanical engineering from the School of Mechanical Engineering, Shanghai Jiao Tong University, Shanghai, China, in 2023. He is currently an Assistant Professor with the Harbin Institute of Technology Shenzhen, Shenzhen, China. His research interests include human–computer interaction, brain functional response, gaze estimation, and their practical applications in medical rehabilitation.
@@
Honghai Liu (Fellow, IEEE) received the Ph.D. degree in robotics from King’s College London, London, U.K., in 2003. He is currently a Professor with the School of Mechanical Engineering and Automation, Harbin Institute of Technology Shenzhen, Shenzhen, China, and with the School of Computing, Portsmouth, U.K. His research interests include human –machine interaction, computer vision, machine learning and their clinical applications with an emphasis on approaches that could make contribution to the intelligent connection of perception to action using contextual information. Prof. Liu was a member of European Academy of Sciences. He is an Associate Editor of IEEE TRANSACTIONS ON INDUSTRIAL CYBERNETICS and the Coeditors-in-Chief of IEEE TRANSACTIONS ON INDUSTRIAL INFORMATICS.

### v2

@@ IV-D Ablation Study
Effect of method’s modules. We conduct an ablation study starting from a baseline model that was trained on the source domain, with our method’s modules systematically integrated to analyze their effects: Pose Augmentor $M_{a}$, Pseudo-Label Generation module $M_{p}$, Global Transformation module $M_{t}$, and iterative training. Since the generated 3D pseudo-labels approximate the target distribution and guide the global transformation, Pseudo-Label Generation module $M_{p}$ and Global Transformation module $M_{t}$ are evaluated as a unified component.
[…]
TABLE V: The impact of each module in the proposed method. The * indicates that the modules were processed for an extra iteration.
@@
VideoPose3D [23] + $M_{a}$ + ($M_{p}$ & $M_{t}$) 58.6 91.8 64.3
VideoPose3D [23] + $M_{a}$ + ($M_{p}$ & $M_{t}$)* 57.7 92.5 64.6
Different Global Transformations. The global transformation module $M_{t}$ contributes significantly to reducing the domain differences in camera positions and orientations from source to target. It is target-oriented and uses the pseudo-labels to guide the rotation and translation of the 3D poses. We compare three other approaches for global transformation: RR+RT, GR+RT, RR+GT, where RR denotes rotation matrices are randomly initialized and learned in the network, RT denotes translation matrices are randomly initialized and learned in the network, GR denotes the pseudo-label guided rotation and GT denotes the pseudo-label guided translation. Our approach is therefore GR+GT, which performs best as shown in Table VI.
TABLE VI: The impact of different approaches for Global Transformation $M_{t}$.
@@
IV-E Analysis of different Global Transformation solutions
We now analyse the reliability of our proposed cross-domain 3D pose alignment to achieve a global transformation against the less computationally efficient Kabsch algorithm [50, 51].
Given the 3D pose sets of the H3.6M dataset and the pseudo-3D pose sets of the 3DHP dataset, we randomly sample an equal number of poses from the H3.6M dataset to match the sample size of the 3DHP dataset. Based on these two equally sized 3D pose sets, we perform random pairwise alignment between them. This process is repeated five times to ensure robustness, and MPJPE is used to quantify the distance between the two sets.
As shown in Table IX, the MPJPE after global transformation alignment shows an average of 219.1mm which is only slightly higher than that of the Kabsch algorithm at 193.9mm.
TABLE IX: MPJPE ($m$) before and after random pairwise global transformation (GT) alignment performed five times. The percentages show the improvement achieved against the ‘Before GT’ value.
1 2 3 4 5 Average $\downarrow$
Before GT 2096.6 2103.1 2117.1 2093.2 2098.2 2101.6
GT using Kabsch
method [50, 51]
193.0 194.1 193.7 193.3 195.5
193.9
(90.8%)
GT using
proposed method
218.1 219.4 218.7 218.6 220.8
219.1
(89.6%)
IV-F Validation of the generated data
To further validate the influence of the generated pseudo-labels on the model, we conducted experiments to train the pose estimator using varying amounts of both the generated pose data and the available target domain pose data. These experiments were carried out using VideoPose3D[23] as the Pose Estimator, with H3.6M as the source domain and 3DHP as the target. The generated pose data was derived from the source domain, at a sample size that matches that of the original source dataset.
As shown in the light-grey part of Table X, for all different proportions of target data deployed, the performance of the model improved across all measures when the generated data was incorporated. In the darker-grey part of Table X, it can be observed that pose estimation performance again improves consistently across all measures when using all the target data and different proportions of generated pose data, ranging from 0% to 100%. For each trial, the proportion of generated data was randomly sampled.
TABLE X: Training the pose estimator using various percentages of generated pose data and target pose data.
[…]
0% 20% 82.5 86.0 52.7
100% 20% 51.7 93.9 67.6
0% 40% 62.8 91.6 61.6
100% 40% 48.4 94.7 69.3
0% 60% 54.8 93.7 65.7
100% 60% 47.6 94.9 69.7
0% 80% 50.6 94.6 67.9
100% 80% 45.8 95.3 70.7
0% 100% 49.0 94.9 68.9
100% 100% 45.0 95.4 71.0
0% 100% 49.0 94.9 68.9
20% 100% 47.2 95.2 69.7
40% 100% 46.4 95.3 70.1
60% 100% 45.8 95.4 70.5
80% 100% 45.3 95.4 70.9
100% 100% 45.0 95.4 71.0
IV-G Discussion Compared to SOTA Methods
@@
The Dual-Augmentor framework in [20] generates diverse out-of-source distributions using a weak and strong augmentor, while our approach fully utilizes target-domain 2D poses to generate pseudo-3D poses and achieves better cross-domain performance. It is primarily applicable when target data are accessible.
@@ V Conclusion
This paper proposes an iterative framework for domain-adaptive human pose estimation, focusing on aligning the 2D and 3D poses across domains to bridge domain gaps. Leveraging 2D poses from the target domain, we investigate a relatively underexplored strategy for estimating absolute 3D poses as pseudo-labels. These target-domain pseudo-labels provide effective guidance for the Global Transformation module, which incorporates a human-centered coordinate system to explicitly align global coordinates between the source and target domains. This design is both intuitive and effective, achieving significant performance gains without adding extra training parameters. Additionally, we enhance pose augmentation by modifying bone lengths and angles, and we propose to use an adversarial learning architecture based on the target domain’s 2D poses and the source domain’s 3D poses. The above modules are cohesively integrated into a unified framework that effectively enhances cross-domain pose estimation performance.
@@ Acknowledgments
Zhiyong Wang (Member, IEEE) received the Ph.D. degree in mechanical engineering from the School of Mechanical Engineering, Shanghai Jiao Tong University, Shanghai, China, in 2023. He is currently an Assistant Professor with the Harbin Institute of Technology Shenzhen, Shenzhen, China. His research interests include human-computer interaction, brain functional response, gaze estimation, and their practical applications in medical rehabilitation.
@@
Honghai Liu (Fellow, IEEE) received the Ph.D. degree in robotics from King’s College London, London, U.K., in 2003. He is currently a Professor with the School of Mechanical Engineering and Automation, Harbin Institute of Technology Shenzhen, Shenzhen, China, and with the School of Computing, Portsmouth, U.K. His research interests include human-machine interaction, computer vision, machine learning and their clinical applications with an emphasis on approaches that could make contribution to the intelligent connection of perception to action using contextual information. Prof. Liu was a member of European Academy of Sciences. He is an Associate Editor of IEEE TRANSACTIONS ON INDUSTRIAL INFORMATICS and the IEEE TRANSACTIONS ON CYBERNETICS.